# A Dense Subframe-based SLAM Framework with Side-scan Sonar

### v1

@@ III-D Pose Graph Optimization
where and is the number of odometry and loop-closure edges in the graph, respectively. is the odometry covariance that can be decided in a similar way as , and is the covariance of loop-closure measurement that comes together with the estimated relative pose. Note that we omit the constant prior factor as shown in Fig. 4 (centre) here to simplify the equation. We solve for Eq. 15 incrementally using iSAM2 [42], i.e., the whole pose graph is updated iteratively from the previous solution, when there are new measurements extracted from an input side-scan image and added to the graph. In this case we can avoid the increase of drift over a long trajectory causing the linearization to be far off that the wrong local minimal is found.
@@ III-E Landmark Reconstruction
After the pose graph optimization, a global optimized AUV pose trajectory is obtained, which is used to reconstruct a quasi-dense map from the dense correspondences in all side-scan images. Specifically, we use Eq. 9 again, but fix the poses with the optimized ones and only optimize for the landmarks. Note that, as the dense correspondences are very noisy and full of outliers due to the edge issues discussed in Section III-B3, we discard part of the landmarks to be quasi-dense in quantity, via threshold on the plane and range costs, which are helpful indicators for recognizing bad quality landmarks.
@@ IV-A1 Dataset Description and Annotation
All the side-scan sonar data tested in this work are collected by Gothenburg University’s Hugin AUV equipped with EdgeTech side-scan sonar (see Fig. 5). The datasets were collecting approximately pings per second with the AUV speed at m/s. The survey lines either following a lawn-mower or a loop pattern are roughly parallel to one another and have large overlap between each other. The seafloor of the surveyed areas are locally flat with gentle slope in altitude, and they contains lots of trawling marks. Details of the dataset and sonar characteristics can be found in Table I.
@@ IV-A2 Evaluation Metrics
Due to lack of ground truth robot pose trajectory, which is often difficult to obtain in underwater, it is hard to evaluate pose trajectory with metric [43] that is commonly used for on-land or aerial robots. To mitigate such problem, we propose to compute the landmark consistency error (LCE) using annotated keypoint correspondences and the bathymetry mesh from MBES data, considering the fact that a landmark observed from different ping pose should have a unique global position in ideal scenario. In particular, for each pair of keypoint correspondence, we project both keypoints onto the mesh by ray-casting and find their intersections with the mesh [44], respectively. Then the Euclidean distance between both intersected landmarks is used as the error metric, i.e., the smaller this consistency error it is, the more accurate the ping poses they are.
@@
To evaluate the bathymetry reconstruction result, one common way is to compute the heightmap given a fixed resolution, then compute the mean absolute error (MAE) against a reference heightmap of MBES mesh [45][46], while this also requires ground truth pose to guarantee the high precision of MBES data. To address this, we propose to evaluate the reconstructed map line-by-line in the sensor frame coordinate. Specifically, we project the estimated landmarks in the global frame coordinate to each survey line in the sonar sensor frame coordinate, and get them aligned with the raw multi-beam point cloud using known sensor offset. Then the MAE between the heightmaps of estimated landmarks and raw multi-beam point cloud is calculated as the error metric, which is free from the influence of pose.
@@ IV-A4 Implementation Details
For all experiments we use the same parameters. In particular, the size (ping number) of each subframe is , and the iteration number . For dense matching, , , . For robust pose estimation, , , , . For landmark reconstruction, the range and plane thresholds are and . For other parameters that are not discussed in the paper, please refer to our source code.
@@ IV-B1 Pose Trajectory
The accuracy of pose estimation is evaluated on Mission1 dataset with annotated keypoints as ground truth reference. Results are shown in Fig. 6. Overall, our proposed method achieve consistent improvement against both the dead-reckoning and ping-based SLAM baselines, with an average error reducing by up to m (). We also add the subframe-based and ping-based SLAM results generated both using the ground truth annotations for comparison. Interestingly, our subframe-based method with much less loop-closing constraints still slightly outperforms the ping-based baseline, which uses every two-ping measurement to estimate pose as loop-closing constraint. This suggests that our subframe-based method can robustly estimate highly accurate relative pose from noisy and outlier-prone measurements.
We also compute the absolute trajectory error (ATE) [43] to evaluate absolute pose consistency, by treating the best result in Fig. 6 as ‘ground truth’ baseline. Fig. 7 demonstrates the error comparison of the whole trajectory, where our estimated trajectory fluctuates slightly in the beginning, while later pulls down the drifts significantly after k pings. As can be seen in Fig. 8 the zoom-in windows, the estimated trajectory is closer to the ’ground truth’ trajectory, with about improvement over dead-reckoning in root mean squared error (RMSE), see Table II. Interestingly, we can notice that the trajectory error with dead-reckoning does not always grow linearly over time, as Hugin running a lawnmower pattern could cancel out the drift growth obtained from body-fixed velocity and heading errors [49]. Similar effect can be observed in the landmark consistency error with dead-reckoning in Fig. 6.
@@ V Conclusion
The proposed method relies on the assumption that the seafloor is relatively flat with gentle slope, primarily in consideration of canonical image transformation for dense matching, which may limits the applicable scenarios, e.g., regions with complex geometric information such as rocks, ridges, hills, etc. However, we may argue that the proposed method would be a good complementary solution for SLAM using multi-beam sensor, which usually requires the seabed to be geometric feature-rich for data association. But still and all, developing a robust and accurate data association algorithm for side-scan images without flat seafloor assumption is a potential direction for future work.

### v2

@@ III-D Pose Graph Optimization
where and is the number of odometry and loop-closure edges in the graph, respectively. is the odometry covariance that can be decided in a similar way as , and is the covariance of loop-closure measurement that comes together with the estimated relative pose. Note that we omit the constant prior factor as shown in Fig. 5 (centre) here to simplify the equation. We solve for Eq. 9 incrementally using iSAM2 [42], i.e., the whole pose graph is updated iteratively from the previous solution, when there are new measurements extracted from an input side-scan image and added to the graph. This iterative approach prevents drift from accumulating over a long trajectory. As such, wrong local minima caused by linearization far away from the optimal solutions can be avoided.
@@ III-E Landmark Reconstruction
After the pose graph optimization, a globally optimized AUV pose trajectory is obtained, which is used to reconstruct a quasi-dense map from the dense correspondences in all side-scan images. Specifically, we use Eq. 5 again, but fix the poses with the optimized ones and only optimize for the landmarks. Note that, as the dense correspondences are very noisy and full of outliers due to the edge issues discussed in Section III-B3, we discard part of the landmarks to be quasi-dense in quantity, via threshold on the plane and range costs, similar to loop closure edge filtering in Section III-C.
@@ IV-A1 Dataset Description and Annotation
All the side-scan sonar data tested in this work are collected by Gothenburg University’s Hugin AUV equipped with EdgeTech side-scan sonar (see Fig. 6). The datasets were collecting approximately pings per second with the AUV speed at m/s. The survey lines either following a lawn-mower or a loop pattern are roughly parallel to one another and have large overlap between each other. The seafloor of the surveyed areas are locally flat with gentle slope in altitude, and they contain lots of trawling marks. Details of the dataset and sonar characteristics can be found in Table I.
@@ IV-A2 Evaluation Metrics
Due to lack of ground truth robot pose trajectory, which is often difficult to obtain in underwater, it is hard to evaluate pose trajectory with metric [43] that is commonly used for on-land or aerial robots. To mitigate such problem, we propose to compute the landmark consistency error (LCE) using annotated keypoint correspondences and the bathymetry mesh from MBES data, considering the fact that a landmark observed from different ping pose should have a unique global position in ideal scenario. In particular, for each pair of keypoint correspondence, we project both keypoints onto the mesh by ray-casting and find their intersections with the mesh [44], respectively. Then, the Euclidean distance between both intersected landmarks is used as the error metric. The smaller this consistency error, the more accurate the ping poses. Note that this metric requires manually annotated keypoints, which is only available for Mission1.
@@
To evaluate the bathymetry reconstruction result, one common way is to compute the heightmap at a fixed resolution, then compute the mean absolute error (MAE) against a reference heightmap of MBES mesh [45][46]. However, this also requires ground truth pose to guarantee the high precision of MBES data. To address this, we propose to evaluate the reconstructed map line-by-line in the sensor frame coordinate. Specifically, we project the estimated landmarks in the global frame coordinate to each survey line in the sonar sensor frame coordinate, and align them with the raw multi-beam point cloud using known sensor offset. Then the MAE between the heightmaps of estimated landmarks and raw multi-beam point cloud is calculated as the error metric, which is free from the influence of pose.
@@ IV-A4 Implementation Details
For all experiments we use the same parameters. In particular, the size (ping number) of each subframe is , and the iteration number . For dense matching, , , . For robust pose estimation, , , , . For landmark reconstruction (loop-closure edge filtering), the range and plane thresholds are and . For other parameters that are not discussed in the paper, please refer to our source code.
@@ IV-B1 Pose Trajectory
As shown in Fig. 7, the accuracy of pose estimation is evaluated on Mission1 dataset by treating the results of running SLAM with annotated keypoints as ground truth references (white and orange). Overall, our proposed method (green) achieve consistent improvement against both the dead-reckoning (red) and ping-based SLAM (blue) baselines, with an average error reduction by up to m (). When comparing only between the ground truth references, interestingly, our subframe-based method (orange) with much less loop-closing constraints still slightly outperforms the ping-based baseline (white), which uses every two-ping measurement to estimate pose as loop-closing constraint. This suggests that our subframe-based method can robustly estimate highly accurate relative pose from noisy and outlier-prone measurements.
@@
We then compute the absolute trajectory error (ATE) [43] to evaluate absolute pose consistency, by treating the trajectory generated by running our proposed SLAM with annotated keypoints (best performance in Fig. 7) as “ground truth” trajectory. Fig. 8 demonstrates the error comparison of the whole trajectory. Although our estimated trajectory fluctuates slightly in the beginning, the ATE is reduced significantly after k pings. As can be seen in Fig. 9 the zoom-in windows, the estimated trajectory is closer to the “ground truth” trajectory, with about improvement over dead-reckoning in root mean squared error (RMSE), see Table II. Interestingly, we can notice that the trajectory error with dead-reckoning does not always grow linearly over time, as Hugin running a lawnmower pattern could cancel out the drift growth obtained from body-fixed velocity and heading errors [49]. Similar effect can be observed in the landmark consistency error with dead-reckoning in Fig. 7.
@@ IV-C2 Iterative Data Association with Optimized Poses
IV-C3 Method Feasibility against DR Accuracy
The proposed methods relies on reasonable odometry solutions (i.e., the Dead-reckoning system in this paper) to work properly. To evaluate the limit of our proposed method against the DR accuracy, we corrupt the Dead-reckoning trajectory by additionally adding different level of Gaussian noise to the heading (yaw) of the vehicle, to simulate larger accumulative drift in the DR estimates, as demonstrated in Fig. 14. With the corrupted DR trajectory as input, we run our proposed method to get an optimized trajectory and compute the ATE against the ”ground truth” reference trajectory, same as in Section IV-B1.
As shown in Fig. 15, our method is able to work with a DR solution with up to m absolute trajectory error, reducing drift by % to m. Nevertheless, it fails to work properly after adding higher level of noises (after rad). The results become worse, mainly due to adding false loop-closure edges, which push certain parts of the trajectory away from the true position.
@@ V Conclusion
The proposed method is limited to reasonable DR solutions to work properly. If rather poor odometry is provided, our proposed system would either fail in finding valid loop closure edge (mistaken overlap check), or obtain invalid loop closure edge (bad data association). Another limitation is dependence on the assumption that the seafloor is relatively flat with gentle slope, primarily in consideration of canonical image transformation for dense matching, which may limits the applicable scenarios, e.g., regions with complex geometric information such as rocks, ridges, hills, etc. However, we may argue that the proposed method would be a good complementary solution for SLAM using multi-beam sensor, which usually requires the seabed to be geometric feature-rich for data association. But still and all, developing a robust and accurate data association algorithm for side-scan images without flat seafloor assumption and/or good odometry prior is a potential direction for future work.